P = 520 – 2Q And where marginal costs are expressed via: MC = 100 + 2Q Within a perfectly competitive market calculate the: a. Consumer surplus b. Producer surplus c. Total surplus Now, consider the market changes to a monopoly, calculate the: a. Consumer surplus b. Producer surplus c. Dead weight loss (otherwise known as the loss to economic surplus)
Answer #1

When the market is Perfectly Competitive

Set P=MC

520-2Q = 100+2Q

520-100 = 2Q+2Q

420 = 4Q

Q = 420/4 = 105

P = 520-2*105 = 310

Maximum reservation price = 520

a) CS = 0.5*105*(520-310) = 11025

b) PS = 0

c) TS = CS+PS = 11025

When the market is a Monopoly

MR = 520-4Q

Set MR=MC

520-4Q = 100+2Q

520-100 = 2Q+4Q

420 = 6Q

Q = 420/6 = 70

P = 520-2*70 = 380

MC = 100+2*70 = 240

a) CS = 0.5*70*(520-380) = 4900

b) PS = 70*(380-240) = 9800

c) DWL = 0.5*(105-70)*(380-310) = 1225